OpenStudy (anonymous):

simplify the expression 5+(radical 7)/5-(radical 7)

OpenStudy (anonymous):

Multiply by the conjugate of the denominator. \[(5+\sqrt{7})/(5-\sqrt{7})*(5+\sqrt{7})/(5+\sqrt{7})\]

OpenStudy (anonymous):

you should end up with \[(32+10\sqrt{7})/18 = (16+5\sqrt{7})/9\]